Here's my question...if the graph of the equation y=f(x), where the vertex of the parabola is 0,0 and P:(2,4), if h(x)=f(x)+3, in (x,y) format what would the new coordinates of P be after the shift? And, if k(x)=-f(x), what would the new coordinates of P be after the reflection? Thanks so much!
Answer by stanbon(75887)

if the graph of the equation y=f(x), where the vertex of the parabola is (0,0) and P:(2,4), if h(x)=f(x)+3, in (x,y) format what would the new coordinates of P be after the shift?
That shift will add 3 to the y-values.
(0,0) will go to (0,3); (2,4) will go to (2,7)
------------------------------------------------
And, if k(x)=-f(x), what would the new coordinates of P be after the reflection?
That shift will flip point over around the x-axis:
(0,0) will not move; (2,4) will ot to (2,-4)
=======================
Cheers,
Stan H.
-------------------------------